OpenStudy (anonymous):

For what x-value(s) does sin(x) = 0? Answer -270°, -90°, 90°, and 270° -270°, -90°, 0°, 90°, and 270° -360°, 0°, and 360° -360°, -180°, 0°, 180°, and 360°

OpenStudy (anonymous):

@jkim3592

OpenStudy (anonymous):

For what x-value(s) does sin(x) = -1? Answer -270°, -90°, 90°, and 270° -270°, -90°, 0°, 90°, and 270° -90 and 270° -270° and 90°

OpenStudy (anonymous):

@hartnn @Hero

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

anybody?

OpenStudy (anonymous):

\[\sin x=0=\sin 180 n\] x=180n,where n is an integer

OpenStudy (anonymous):

ok, that is a bit confusing to me :/

OpenStudy (anonymous):

\[\sin x=-1=-\sin 90=\sin \left( -90 \right),\sin \left( 180+90 \right)\]

OpenStudy (anonymous):

x=180n,n=....-2,-1,0,1,2...

OpenStudy (anonymous):

ok but, thats not an option of answers :/

OpenStudy (anonymous):

last one i see in your option. i.e.,x=-2*180=-360 n=-1,x=-1*180=-180 and so on for x=0,1,2

OpenStudy (anonymous):

x=-90 x=270

OpenStudy (anonymous):

correction for n=0,1,2

OpenStudy (anonymous):

oh! lol thanks. um, is that for the first one or the second one though?

OpenStudy (anonymous):

-90,270 for second, -360,-180,0,180,360 for first

OpenStudy (anonymous):

thank you! you are a lifesaver!